volume of box is 24 cubic feet. height is 2 feet and width is 4 feet. what is the length of the box?
Found 2 solutions by unlockmath, lwsshak3:
Answer by unlockmath(1688)

Hello,
Follow this formula WHL=Volume
Plug in the numbers:
2(4)L=24
Divide by 8
L=3 feet
Make sense?
RJ
www.math-unlock.com

Answer by lwsshak3(11628)

volume of box is 24 cubic feet. height is 2 feet and width is 4 feet. what is the length of the box?
*
Volume=length*width*height
24=length*4*2=length*8
length=24/8=3 feet